OpenStudy (anonymous):

Simplify i^{38} i -1 -i 1

OpenStudy (anonymous):

Alright lets do this together.. first what do you think the answer is @Tatsurion

OpenStudy (anonymous):

1 is my guess.

OpenStudy (anonymous):

I believe you are right. You might want to confirm with someone else though :) Although, while I am here you can go to http://aihberliterally.wix.com/aihberliterally to book free math tutoring lessons.

OpenStudy (anonymous):

Have a great day!

OpenStudy (anonymous):

Says it's -1. https://mathway.com/popular-problems/Algebra/201886

OpenStudy (danjs):

there is a simple way to figure those when you keep adding another i power a few times, you will notice a cycle every 4 terms repeat. so you can divide that large number by 4, and figure where in that 4 term sequence the anser is

OpenStudy (danjs):

i = i i^2 = -1 i^3 = i^2*i = -i i^4 = +1 then it repeats again, , like i^5 is just i^4 *i = i, same as i^1

OpenStudy (anonymous):

So it is 1?

OpenStudy (danjs):

since that cycle repeats, you can take all the 4's out of 38, and the remainder will be the term in that 4 term sequence to use 4 goes into 38, 9 times, left over with 2 reaining, must be the same as i^2, the second of the 4 terms

OpenStudy (danjs):

-1

OpenStudy (anonymous):

Oh! Now I get it.

OpenStudy (danjs):

or maybe think of it as, i^4 is just 1 factor out all the i^4 you can , those go to just multiplying by 1, left with the same answer